OpenStudy (butterflyhope):

Help would be appreciated :) Consider the equality x - 1/ x + 1 < 3. When solving this inequality, should the inequality sign change direction? Why or why not? Thanks c:

OpenStudy (campbell_st):

it will change direction because to solve it you need to multiply by the square of the denominator

OpenStudy (calculusfunctions):

The inequality sign should change direction when i). Multiplying or dividing both sides of an inequality by a negative, ii). reading an inequality backwards, or iii). when taking the reciprocal of both sides iff both sides are of the same sign (positive or negative).

OpenStudy (calculusfunctions):

Sorry, my internet is having issues.

OpenStudy (campbell_st):

this is how I would solve this inequality. start with x cannot equal - 1 mulltiply both sides by the square of the denominator \[\frac{x -1}{x + 1} \times (x + 1)^2 < 3\times (x + 1)^2\] which becomes \[x^2 - 1 < 3x^2 + 6x + 3\] or \[-2x^2 - 6x - 4 < 0\] multiply by -1/2 will reverse the inequality \[x^2 + 3x + 2 > 0\] \[(x + 2)(x + 1) > 0\] so the possible solutions are x = -2 and x = -1 now test the solution  hope it helps